OpenStudy (anonymous):

Simplify the square root 2,475. 9 the square root of 30 11 the square root of 15 15 the square root of 11 45 the square root of 11

OpenStudy (anonymous):

Where should I start?

OpenStudy (anonymous):

We have : \[\Large 2,475=\frac{2475}{1000}=\frac{25\times 9\times 11}{100\times 10}\] So : \[\sqrt{2,475}=\frac{5\times3\sqrt11}{10\sqrt{10}}=\frac{15\sqrt{11}\sqrt{10}}{10\sqrt{10}\sqrt{10}} \\=\frac{15\sqrt{110}}{100}=\frac{3\sqrt{110}}{20} \]